OpenStudy (anonymous):

find the current through the resistor and the charge on the positive plate of the capacitor after 17 seconds. 13Voltage 4ohms (resistor) 6F (capacitor)

OpenStudy (anonymous):

c=q/v, and I=v/r,

OpenStudy (anonymous):

\[V=IR\] V=Voltage I=current R=Resistance